# A Good Foundation is Worth Many Labels: Label-Efficient Panoptic Segmentation

### v1

@@ Abstract
I Introduction
@@
In this work, we employ this paradigm shift to panoptic segmentation to substantially reduce the number of annotated images required for training. In particular, we propose a novel approach for PAnoptic SegmenTation with fEw Labels (PASTEL) that is based on three main pillars. First, PASTEL exploits descriptive DINOv2 [15] image features to perform both semantic segmentation and object boundary detection. Second, PASTEL comprises a novel panoptic fusion module to merge these predictions in a post-processing step that not only produces the panoptic output but also refines the semantic map. To separate instances, PASTEL utilizes recursive two-way normalized cut [16]. Finally, PASTEL bootstraps selectively sampled unlabeled images to further improve its performance via self-training. In extensive experiments, we demonstrate that PASTEL creates high-quality panoptic predictions from as few as 10 labeled images on Cityscapes [4], Pascal VOC [17], and PhenoBench [18]. Notably, PASTEL can hence be trained with labels produced by a single annotator in 1 days [4] while outperforming previous label-efficient methods that require five to ten times as much data. We further show that the predictions of PASTEL can be used as pseudo-labels to train densely supervised models, i.e., rendering them label-efficient. To encourage future research toward label-efficient segmentation with foundation models, we release our code at http://pastel.cs.uni-freiburg.de.
@@ II Related Work
Label-Efficient Image Segmentation. Classical deep image segmentation methods require a large amount of annotated training data [1, 2, 3]. Therefore, many recent works employ different strategies of weak supervision [26] to reduce the labeling cost. For instance, unsupervised semantic segmentation is commonly addressed using contrastive learning techniques [10, 27] to find similar clusters in the feature space. Recent methods have leveraged descriptive image representations from large-scale task-agnostic pretraining [25] for both semantic [12] and instance segmentation [11]. With respect to the more challenging task of panoptic segmentation, CoDEPS [28] distills knowledge from a labeled source domain to a new unlabeled target domain. Sparse annotations offer an intermediate approach with limited pixel-based supervision generated by an inexpensive labeling scheme, e.g., point annotations [5, 6]. Semi-supervised training bootstraps a few densely annotated examples with a large set of unlabeled images. Commonly, semi-supervised semantic segmentation is based on auxiliary tasks [7], self-training [8], or uncertainty estimation [9]. Finally, visual foundation models [15, 24, 25] enable label-efficient densely supervised segmentation due to their semantically rich image representations, e.g., SPINO [13] performs panoptic segmentation from very few annotated images. Inspired by these recent insights, we exploit DINOv2 [15] features to perform panoptic segmentation with a minimum number of labels. Unlike most prior label-efficient techniques, in this work we aim to enable panoptic segmentation from only as few labeled images as a single annotator can produce within a reasonable time frame, thus facilitating deployment in custom domains.
@@ III-A Model Architecture and Training
The key insight of our PASTEL is to exploit the semantically rich image features from a foundation model to perform label-efficient segmentation and instance delineation.
Network Design. As illustrated in Fig. 2, we design our network according to the multi-task paradigm with a shared backbone. Inspired by Point2Mask [5], we separately perform pixel-based semantic segmentation and object boundary detection. In detail, we employ the pretrained ViT-B/14 variant of DINOv2 [15] as the frozen backbone. In the -class segmentation head, we first upsample the patch-wise features of DINOv2 to the input image size, i.e., -upsampling. We then feed the output to four 1 convolution layers of feature sizes , , , and . In the object boundary head, we operate on a smaller feature map using a -upsampling layer, again followed by four 1 convolution layers of output sizes , , , and . We frame the boundary detection task as binary classification with labels and denoting boundary and background pixels, respectively. During test-time, the output of both heads is merged by our novel panoptic fusion module as detailed in Sec. III-B.
@@ III-B Panoptic Fusion Module
to prevent sharp borders within the merged prediction that can result in visual artifacts. The parameter defines the extent of the overlap, e.g., indicates that half of an image crop is overlapped by another crop. We depict this method in Fig. 3 for scale and overlap , yielding nine image crops. We upsample each crop to the input image size using bilinear interpolation and feed them through the respective head. Then, we downsample the generated feature maps to and place them in a combined feature map at the position corresponding to the input image crop. We repeat this procedure for each scale and average features of overlapping pixels. Finally, we merge the features from multiple scales using the mean value per pixel.
Panoptic Fusion and Refinement. We visualize the individual steps of our proposed panoptic fusion module in Fig. 2, starting with the previously described multi-scale fusion (1). First, we compute an affinity matrix from the predicted boundary map (2) to be used in the instance delineation of step (9), detailed in the next paragraph. In step (3), we obtain the binary boundary map after thresholding the class probabilities. We further denoise by removing small boundaries. Next, we extract boundaries between any two “stuff” and “thing” classes and add them to (4). This enables us to find disconnected segments in the predicted semantic map . In detail, we perform connected component analysis (CCA) on , followed by majority voting over the semantic classes present in a segment (5). In Fig. 2, this changes the burgundy colored pixels in the left vehicles to the car class. In step (6), we again perform CCA but using the semantic predictions, i.e., we obtain a segment for each area in the image whose neighbors belong to a different semantic category. In the following, we separate these segments into “stuff” and “thing” segments. First, we iterate over the “thing” segments (7) and set the semantic label to the ignore class if the segment size is below a threshold, the boundary head does not predict a boundary for any of the segments’ pixels, or the segment is fully surrounded by another thing class. While the second filter is inspired by ensemble learning, the third filter targets infeasible objects flying in the scene. In Fig. 2, these filters remove the red pixels in the rear window of the center vehicle as well as the smaller car segments. Then, we iterate over the “stuff” segments (8) and apply the same filters except for the boundary-based removal. In step (9), we fuse the semantic prediction with the detected object boundaries to obtain a panoptic map. Finally, in step (10), we propagate labels from the nearest neighbor to fill all previously created holes, i.e., pixels that were assigned the ignore label. Note that although we only visualize car and road classes in Fig. 2, the final panoptic segmentation map contains all valid categories.
@@
In step (2), we first downsample the boundary map to size . Then, we compute a sparse affinity matrix from the estimated soft boundary map, i.e., before thresholding, as follows: We create a sparse distance matrix , where for each pixel we draw a line with the Bresemham algorithm to all neighbors within a specified radius. The distance between pixels and is defined as
where is the predicted soft boundary map. We convert the distances to affinities by taking the negative exponential:
@@
where is a diagonal matrix with diagonal elements . The solution is a continuous bipartition of the segment. If the cost of a cut, represented by , is less than a threshold and a stability criterion is fulfilled, we apply the cut. Otherwise, we stop the recursion. The stability criterion [16] measures the degree of smoothness in . Formally, if the ratio between the minimum and the maximum value of a histogram of is less than a threshold, the criterion holds. Thus, we do not cut if there is high uncertainty in the solution . To finally apply the cut, we search for the splitting point that minimizes the NCut cost to bipartition into two segments. We then recursively apply this procedure to both segments to find additional instances. If the size of a segment is below a threshold, we remove this segment and set its semantic label to the ignore class. In contrast to CCA, employing NCut allows for handling small gaps and noise present in the boundary map.
@@ IV Experimental Evaluation
We demonstrate that our PASTEL method outperforms previous label-efficient segmentation techniques while requiring significantly fewer annotations. In extensive ablation studies, we analyze the various design choices.
IV-A Datasets
We present results on three diverse datasets: First, the Cityscapes dataset [4] provides RGB images and high-quality panoptic annotations for urban driving. For evaluation, we report results on the val split using the official evaluation protocol with 19 classes. Note that the entire train split comprises 2,975 images. Second, the Pascal VOC 2012 dataset [17] was originally proposed as an object detection benchmark and has been substantially extended by SBD [31]. Concerning panoptic segmentation, the dataset comprises 20 “thing” classes and a single “stuff” class representing the background. We provide results on the val split. For training, we select images from the train split containing a total of 10,582 images. Finally, the PhenoBench dataset [18] comprises several segmentation tasks for the agricultural domain. We apply our method to the leaf instance segmentation challenge. We select our training samples from the train split, which comprises 1,407 images, and provide results for both val and test splits. In stark contrast to autonomous driving, agricultural robotics lacks large-scale datasets underlining the importance of highly label-efficient approaches.
| Method | Backbone | Supervision | mIoU | PQ |
| Mask2Former [2] | Swin-L | 82.9 | 66.6 | |
| Mask2Former‡ | Swin-L | 50.7 | 29.2 | |
| PiCIE† [10] | ResNet-18 | 13.8 | – | |
| STEGO† [12] | DINO | 38.0 | – | |
| ST++ [8] | ResNet-50 | + | 61.4 | – |
| ST++ [8] | ResNet-50 | 55.1 | – | |
| Hoyer et al. [7] | ResNet-101 | 62.1 | – | |
| Hoyer et al.‡ | DINOv2 ViT-B | 46.4 | – | |
| ST++‡ | DINOv2 ViT-B | 53.0 | – | |
| PanopticDeepLab† [1] | DINOv2 ViT-B | 49.4 | 20.6 | |
| SPINO [13] | DINOv2 ViT-B | 61.2 | 36.5 | |
| PASTEL (ours) | DINOv2 ViT-L | 75.5 | 50.7 | |
| PASTEL (ours) | DINOv2 ViT-S | 64.2 | 41.0 | |
| PASTEL (ours) | DINOv2 ViT-B | 63.3 | 41.3 | |
| PASTEL (ours) | DINOv2 ViT-B | 64.8 | 42.4 |
| Method | Backbone | Supervision | mIoU | PQ |
| Panoptic FCN [6] | ResNet-50 | 80.2 | 67.9 | |
| MaskContrast [27] | ResNet-50 | 35.0 | – | |
| MaskDistill [32] | ResNet-50 | 48.9 | – | |
| U2PL [9] | ResNet-101 | + | 68.0 | – |
| ST++ [8] | ResNet-50 | + | 65.2 | – |
| U2PL [9] | ResNet-101 | 45.8 | – | |
| ST++ [8] | ResNet-50 | 50.7 | – | |
| ST++‡ | DINOv2 ViT-B | 52.8 | – | |
| PASTEL (ours) | DINOv2 ViT-L | 71.1 | 47.3 | |
| PASTEL (ours) | DINOv2 ViT-B | 60.6 | 37.0 | |
| PASTEL (ours) | DINOv2 ViT-B | 62.5 | 39.5 |
@@ IV-A Datasets
| Method | Backbone | Supervision | PQ (val) | PQ (test) |
| Mask R-CNN [33] | ResNet-50 | 61.5 | 59.7 | |
| Mask R-CNN‡ | ResNet-50 | 41.5 | 38.2 | |
| PASTEL (ours) | DINOv2 ViT-B | 51.7 | 49.0 |
The subscript denotes the number of labeled training images. ‡: Baseline trained by us.
@@ IV-B Panoptic Segmentation Results
We compare PASTEL to other label-efficient semantic and panoptic segmentation methods. We further showcase how to render a state-of-the-art segmentation model label-efficient by using PASTEL as a plugin.
@@
In Tab. I, we report results for the Cityscapes dataset. We first address the main target use case by allowing only ten annotated images for training. Our PASTEL achieves remarkable mIoU and PQ corresponding to an increase of mIoU and PQ to the recent SPINO [13]. Notably, the state-of-the-art works Mask2Former [2] and PanopticDeepLab [1] adapted with a DINOv2 [15] backbone perform significantly worse revealing the need for specifically designed methods for extreme label efficiency. We further evaluate the work from Hoyer et al. [7] and ST++ [8] but replace the backbones with DINOv2 ViT-B to eliminate their impact. Besides the lack of instance predictions, this results in and mIoU compared to PASTEL. After employing our proposed self-training strategy, the improvement over the concurrent SPINO [13] further increases to mIoU and PQ, yielding also higher metrics than the semi-supervised ST++ [8] that is trained with more labels. For the third case, we use PASTEL with a DINOv2 ViT-S backbone (86M param.) and compare it with the ResNet-101-based [34] Hoyer et al. [7] (45M param.) when training on the same 100 annotations. Our approach yields mIoU plus instance predictions. Finally, we show the potential of PASTEL with a DINOv2 ViT-L backbone that expands the increase to mIoU. Notably, this reduces the gap to the fully supervised Mask2Fomer [2] to mIoU and PQ while using only of the labels.
In Tab. II, we repeat similar experiments on the Pascal VOC dataset. When using only 20 annotated images, PASTEL yields mIoU and PQ, outperforming previous densely supervised methods with limited samples. Compared to ST++ [8] with a DINOv2 ViT-B backbone, the metrics represent mIoU when trained with the same images. Similar to Cityscapes, self-training further increases the performance of PASTEL. For the third setup, PASTEL achieves an increase of mIoU versus the supervision-only baseline reported in ST++ [8] when using the same number of annotations. On Pascal VOC, we can reduce the gap to the fully supervised Panoptic FCN [6] to mIoU and PQ while using only of the labels.
In Tab. III, we report results for the PhenoBench leaf instance segmentation task. Separating leaves is essential to assess the growth stage of crops and to detect diseases. However, annotating a conventional training set with hundreds of images is infeasible. Thus, we demonstrate that PASTEL significantly outperforms the best-performing baseline Mask R-CNN [33] from the dataset’s benchmark, when only of the labeled images are available.
@@
Usage as Pseudo-Label Generator. In this experiment, we leverage the panoptic predictions of PASTEL as pseudo-labels to train a densely supervised panoptic segmentation model. In detail, we train Mask2Former [2] with a ResNet-50 [34] backbone using the official code for three different settings. First, we use the ground truth annotations of the train split of Cityscapes. Second, we generate panoptic pseudo-labels for the same data using PASTEL with the setting. Finally, we add pseudo-labels on the train_extra split showing how to leverage large unlabeled datasets with our method. On all pseudo-labels, we mask the static hood of the ego vehicle following previous works [35, 13]. We report the performance on val data in Tab. IV. Note that the numbers from the authors are slightly greater than our reproduced results, mIoU and PQ [2], but do not include SQ and RQ metrics. Importantly, the panoptic metrics (PQ, SQ, RQ) with train pseudo-labels exceed the results obtained directly with PASTEL, i.e., training Mask2Former further bootstraps the panoptic segmentation performance without increasing the utilized number of human annotations. When adding the train_extra pseudo-labels, the panoptic segmentation scores can be further improved, achieving PQ compared to the results of PASTEL. In summary, this experiment not only demonstrates that PASTEL can serve as a plugin rendering existing densely supervised segmentation models label-efficient but also enables real-time inference [2].

### v2

@@ Abstract
Index Terms:
Semantic Scene Understanding; Deep Learning Methods; Computer Vision for TransportationI Introduction
@@
In this work, we employ this paradigm shift to panoptic segmentation to substantially reduce the number of annotated images required for training. In particular, we propose a novel approach for PAnoptic SegmenTation with fEw Labels (PASTEL) and illustrate the key idea in Fig. 1. First, the vision foundation model DINOv2 [15] and a small set of densely annotated images form the basis of PASTEL. Second, the descriptive image features of DINOv2 [15] allow for highly label-efficient training of two lightweight heads for semantic segmentation and object boundary estimation. Third, at inference time, a novel panoptic fusion module then merges the task-specific predictions and further refines their quality. Finally, PASTEL bootstraps selectively sampled unlabeled images for an additional performance boost via self-training. In extensive experiments, we demonstrate that PASTEL creates high-quality panoptic predictions from as few as 10 labeled images on Cityscapes [4], Pascal VOC [16], and PhenoBench [17]. Notably, PASTEL can hence be trained with labels produced by a single annotator in 1 days [4] while outperforming previous label-efficient methods that require five to ten times as much data. We further show that the predictions of PASTEL can be used as pseudo-labels to train densely supervised models, i.e., rendering them label-efficient. To encourage future research, we release our code at http://pastel.cs.uni-freiburg.de.
@@ II Related Work
Label-Efficient Image Segmentation. Classical deep image segmentation methods require a large amount of annotated training data [1, 2, 3]. Therefore, many recent works employ different strategies of weak supervision [25] to reduce the labeling cost. For instance, unsupervised semantic segmentation is commonly addressed using contrastive learning techniques [10, 26] to find similar clusters in the feature space. Recent methods have leveraged descriptive image representations from large-scale task-agnostic pretraining [24] for both semantic [12] and instance segmentation [11]. With respect to the more challenging task of panoptic segmentation, CoDEPS [27] distills knowledge from a labeled source domain to a new unlabeled target domain. Sparse annotations offer an intermediate approach with limited pixel-based supervision generated by an inexpensive labeling scheme, e.g., point annotations [5, 6]. Semi-supervised training bootstraps a few densely annotated examples with a large set of unlabeled images and is commonly based on auxiliary tasks [7], self-training [8], or uncertainty estimation [9]. In this work, we add to the promising line of research that exploits visual foundation models for label-efficient dense supervision. The concurrent SPINO [13] approach combines a DINOv2 [15] backbone with separate heads for semantic segmentation and object boundary estimation. Inspired by these recent insights, we follow a similar design scheme but exploit further synergies between semantic classes and leverage unlabeled images via self-training. Importantly, unlike most prior label-efficient techniques, in this work we aim to enable panoptic segmentation from only as few labeled images as a single annotator can produce within a reasonable time frame, thus facilitating deployment in custom domains.
@@ III-A Model Architecture and Training
The key insight of our PASTEL is to exploit the semantically rich image features from a foundation model to enable label-efficient segmentation and instance delineation.
Network Design. As illustrated in Fig. 2, we design our network according to the multi-task paradigm with a shared backbone. Inspired by the approaches Point2Mask [5] and SPINO [13], we separately perform pixel-based semantic segmentation and object boundary detection while using a shared backbone. In detail, we employ the pretrained ViT-B/14 variant of DINOv2 [15] as the frozen backbone. In the -class segmentation head, we first upsample the patch-wise features of DINOv2 to the input image size, i.e., -upsampling. We then feed the output to four 1 convolution layers of feature sizes , , , and . In the object boundary head, we operate on a smaller feature map using a -upsampling layer, again followed by four 1 convolution layers of output sizes , , , and . We frame the boundary detection task as binary classification with labels and denoting boundary and background pixels, respectively. During test-time, the output of both heads is merged by our novel panoptic fusion module as detailed in Sec. III-B.
@@ III-B Panoptic Fusion Module
Unlike non-overlapping ensembles [13], our approach prevents sharp borders within the merged prediction that can result in visual artifacts. The parameter defines the extent of the overlap, e.g., indicates that half of an image crop is overlapped by another crop. We depict this method in Fig. 3 for scale and overlap , yielding nine image crops. We upsample each crop to the input image size using bilinear interpolation and feed them through the respective head. Then, we downsample the generated feature maps to and place them in a combined feature map at the position corresponding to the input image crop. We repeat this procedure for each scale and average features of overlapping pixels. Finally, we merge the features from multiple scales using the mean value per pixel.
Panoptic Fusion and Refinement. We visualize the individual steps of our proposed panoptic fusion module in Fig. 2, starting with the previously described multi-scale prediction (1). First, we compute an affinity matrix from the predicted soft boundary (2) to be used for instance delineation (9). We detail these steps in the next paragraph. In step (3), we obtain the binary boundary after thresholding the class probabilities of each element with :
We further denoise by removing small boundaries. Next, we extract boundaries between any two “stuff” and “thing” classes (4) and add them to . This enables us to find disconnected segments in the predicted semantic map . In detail, in step (5) we perform connected component analysis (CCA) on , followed by majority voting to update the pixel-wise predicted semantic classes of a segment :
In Fig. 2, this changes the burgundy colored pixels in the left vehicles to the car class. In step (6), we again perform CCA but use the semantic predictions, i.e., we obtain a segment for each area in the image whose neighbors belong to a different semantic category. In the following, we separate these segments into “stuff” and “thing” segments. First, we iterate over the “thing” segments (7) and set the semantic label to the ignore class if the segment size is below a threshold, the boundary head does not predict a boundary for any of the segments’ pixels, or the segment is fully surrounded by another thing class. While the second filter is inspired by ensemble learning, the third filter targets infeasible objects flying in the scene. In Fig. 2, these filters remove the red pixels in the rear window of the center vehicle as well as the smaller car segments. Then, we iterate over the “stuff” segments (8) and apply the same filters except for the boundary-based removal. In step (9), we fuse the semantic prediction with the detected object boundaries to obtain a panoptic map. Finally, in step (10), we propagate labels from the nearest neighbor to fill all previously created holes, i.e., pixels that were assigned the ignore label. Note that although we only visualize car and road classes in Fig. 2, the final panoptic segmentation map contains all valid categories.
@@
In step (2), we first downsample the soft boundary map to size . Then, we compute a sparse affinity matrix based on the distance matrix with distances between pixels and defined as:
where where the operator is provided by the Bresenham algorithm. We convert the distances to affinities by taking the negative exponential:
@@
where is a diagonal matrix with diagonal elements . The solution is a continuous bipartition of the segment. If the cost of a cut, represented by , is less than a threshold and a stability criterion is fulfilled, we apply the cut. Otherwise, we stop the recursion. The stability criterion [29] measures the degree of smoothness in . Formally, if the ratio between the minimum and the maximum value of a histogram of is less than a threshold, the criterion holds. Thus, we do not cut if there is high uncertainty in the solution . To finally apply the cut, we search for the splitting point that minimizes the NCut cost to bipartition into two segments. We then recursively apply this procedure to both segments to find additional instances. If the size of a segment is below a threshold, we remove this segment and set its semantic label to the ignore class. While using CCA [13] fails for non-closed object boundaries, NCut is robust to small gaps and noise present in the boundary map.
@@ IV Experimental Evaluation
We demonstrate that our PASTEL method outperforms previous label-efficient segmentation techniques while requiring significantly fewer annotations. We further showcase that using PASTEL as a plugin can render state-of-the-art segmentation models label-efficient. In extensive ablation studies, we analyze the various design choices.
IV-A Datasets and Implementation Details
We present results on three diverse datasets: First, the Cityscapes dataset [4] provides RGB images and high-quality panoptic annotations with 19 classes for urban driving. Second, the Pascal VOC 2012 dataset [16] was originally proposed as an object detection benchmark and has been substantially extended by SBD [31]. Concerning panoptic segmentation, the dataset comprises 20 “thing” classes and a single “stuff” class representing the background. Finally, the PhenoBench dataset [17] comprises several segmentation tasks for the agricultural domain. We apply our method to the leaf instance segmentation challenge. In stark contrast to autonomous driving, agricultural robotics lacks large-scale datasets underlining the importance of highly label-efficient approaches. In our experiments, we select images from the train split of the respective dataset and report results on the val split. On PhenoBench, we provide further metrics for the test split. If not noted otherwise, we train both heads for 150 epochs on an Nvidia RTX A6000 GPU taking and for semantic segmentation and object boundary estimation, respectively. The inference time with our default settings is approx. per image. We discuss real-time deployment in the last paragraph of Sec. IV-B.
| Method | Backbone | Supervision | mIoU | PQ | |
| 1a) | Mask2Former [2] | Swin-L | 82.9 | 66.6 | |
| 2a) | PiCIE† [10] | ResNet-18 | 13.8 | – | |
| 2b) | STEGO† [12] | DINO | 38.0 | – | |
| 3a) | ST++ [8] | ResNet-50 | + | 61.4 | – |
| 4a) | ST++ [8] | ResNet-50 | 55.1 | – | |
| 4b) | Hoyer et al. [7] | ResNet-101 | 62.1 | – | |
| 4c) | Hoyer et al.‡ | DINOv2 ViT-B | 46.4 | – | |
| 4d) | ST++‡ | DINOv2 ViT-B | 53.0 | – | |
| 4e) | PanopticDeepLab† [1] | DINOv2 ViT-B | 49.4 | 20.6 | |
| 4f) | Mask2Former‡ | Swin-L | 50.7 | 29.2 | |
| 4g) | SPINO [13] | DINOv2 ViT-B | 61.2 | 36.5 | |
| 5a) | PASTEL (ours) | DINOv2 ViT-L | 75.5 | 50.7 | |
| 5b) | PASTEL (ours) | DINOv2 ViT-S | 64.2 | 41.0 | |
| 5c) | PASTEL (ours) | DINOv2 ViT-B | 63.3 | 41.3 | |
| 5d) | PASTEL (ours) | DINOv2 ViT-B | 64.8 | 42.4 | |
| Method | Backbone | Supervision | mIoU | PQ | |
| 1a) | Panoptic FCN [6] | ResNet-50 | 80.2 | 67.9 | |
| 2a) | MaskContrast [26] | ResNet-50 | 35.0 | – | |
| 2b) | MaskDistill [32] | ResNet-50 | 48.9 | – | |
| 3a) | U2PL [9] | ResNet-101 | + | 68.0 | – |
| 3c) | ST++ [8] | ResNet-50 | + | 65.2 | – |
| 4a) | U2PL [9] | ResNet-101 | 45.8 | – | |
| 4b) | ST++ [8] | ResNet-50 | 50.7 | – | |
| 4c) | ST++‡ | DINOv2 ViT-B | 52.8 | – | |
| 5a) | PASTEL (ours) | DINOv2 ViT-L | 71.1 | 47.3 | |
| 5b) | PASTEL (ours) | DINOv2 ViT-B | 60.6 | 37.0 | |
| 5c) | PASTEL (ours) | DINOv2 ViT-B | 62.5 | 39.5 | |
@@ IV-A Datasets and Implementation Details
| Method | Backbone | Supervision | PQ (val) | PQ (test) | |
| 1a) | Mask R-CNN [33] | ResNet-50 | 61.5 | 59.7 | |
| 2a) | Mask R-CNN‡ | ResNet-50 | 41.5 | 38.2 | |
| 2b) | PASTEL (ours) | DINOv2 ViT-B | 51.7 | 49.0 | |
In 2a) and 2b), we used 15 images for training. ‡: Baseline trained by us.
@@ IV-B Panoptic Segmentation Results
For evaluation of PASTEL and other baseline methods, we report mIoU and PQ metrics [34]. While the mIoU only refers to semantic segmentation, the PQ measures the quality of performing panoptic segmentation.
@@
In Tab. I, we report results for the Cityscapes dataset. We first address the main target use case by allowing only ten annotated images for training. Our PASTEL (5c) achieves remarkable mIoU and PQ corresponding to an increase of mIoU and PQ to the recent SPINO [13] (4g). Notably, the state-of-the-art works Mask2Former [2] (4f) and PanopticDeepLab [1] adapted with a DINOv2 [15] backbone (4e) perform significantly worse revealing the need for specifically designed methods for extreme label efficiency. We further evaluate the work from Hoyer et al. [7] (4c) and ST++ [8] (4d) but replace the backbones with DINOv2 ViT-B to eliminate their impact. Besides the lack of instance predictions, this results in and mIoU compared to PASTEL. After employing our proposed self-training strategy (5d), the improvement over the concurrent SPINO [13] further increases to mIoU and PQ, also yielding higher metrics than the semi-supervised ST++ [8] (3a) that is trained with more labels. For the third case, we use PASTEL with a DINOv2 ViT-S backbone (86M param.) (5b) and compare it with the ResNet-101-based [35] Hoyer et al. [7] (45M param.) (4b) when training on the same 100 annotations. Our approach yields mIoU plus instance predictions. Finally, we show the potential of PASTEL with a DINOv2 ViT-L backbone (5a) that expands the increase to mIoU. Notably, this reduces the gap to the fully supervised Mask2Fomer [2] (1a) to mIoU and PQ while using of the labels.
In Tab. II, we repeat similar experiments on the Pascal VOC dataset. When using only 20 annotated images, PASTEL (5b) yields mIoU and PQ, outperforming previous densely supervised methods with limited samples. Compared to ST++ [8] with a DINOv2 ViT-B backbone (4c), the metrics represent mIoU when trained with the same images. Similar to Cityscapes, self-training further increases the performance of PASTEL (5c). For the third setup, PASTEL (5a) achieves an increase of mIoU versus the supervision-only baseline reported in ST++ [8] (4b) when using the same number of annotations. On Pascal VOC, we can reduce the gap to the fully supervised Panoptic FCN [6] (1a) to mIoU and PQ while using only of the labels.
In Tab. III, we report results for the PhenoBench leaf instance segmentation task. Separating leaves is essential to assess the growth stage of crops and to detect diseases. However, annotating a conventional training set with hundreds of images is infeasible. Thus, we demonstrate that PASTEL (2b) significantly outperforms the best-performing baseline Mask R-CNN [33] from the dataset’s benchmark, when only of the labeled images are available (2a).
@@
Usage as Pseudo-Label Generator. In this experiment, we leverage the panoptic predictions of PASTEL as pseudo-labels to train a densely supervised panoptic segmentation model. In detail, we train Mask2Former [2] with a ResNet-50 [35] backbone using the official code for three different settings. First, we use the ground truth annotations of the train split of Cityscapes. Second, we generate panoptic pseudo-labels for the same data using PASTEL with the setting. Finally, we add pseudo-labels on the train_extra split showing how to leverage large unlabeled datasets with our method. On all pseudo-labels, we mask the static hood of the ego vehicle following previous works [36, 13]. We report the performance on val data in Tab. IV. Note that the numbers from the authors are slightly greater than our reproduced results, mIoU and PQ [2], but do not include SQ and RQ metrics. Importantly, the panoptic metrics with train pseudo-labels exceed the results obtained directly with PASTEL, i.e., training Mask2Former further bootstraps the panoptic segmentation performance without increasing the utilized number of human annotations. When adding the train_extra pseudo-labels, the panoptic segmentation scores can be further improved, achieving PQ compared to the results of PASTEL. In summary, this experiment not only demonstrates that PASTEL can serve as a plugin rendering existing densely supervised segmentation models label-efficient but also enables real-time inference [2].